# Modular Jump Gaussian Processes

### v1

@@ 3.2 An Illustration
Now consider applying Alg. 1 independently on each $x\in\mathcal{X}$, e.g., testing on a $101\times 101$ grid, extending our running “Phantom” example. The right panel of Figure 4 uses color to indicate the predictive means obtained at each $x$ on the last line of the algorithm. Compared to the truth [Figure 3] the fit looks good, visually speaking, but not perfect. We show in Appendix B.1 that this fit is more accurate, in terms of RMSE, than an ordinary GP fit. Here we aim to illustrate the optimal choice of neighborhood size, $\hat{n}(x)$, in two ways. We begin with focus on two particular predictive locations: $(-0.32,0.17)$ and $(0.39,0.34)$, far from and near to the manifold of discontinuity, respectively. Observe how the optimal neighborhood size (OLAGP) adjusts as appropriate ($n=43$ far from the manifold; $n=14$ near).
Size OLAGP Modular JGP
12–20 4548 3914
21–40 2838 2853
41–60 745 933
61–80 414 697
81–100 319 298
101–120 139 396
121–140 511 544
141–160 687 566
Figure 5: Distribution of OLAGP and Modular JGP [Section 4] estimated neighborhood sizes.
Now consider how all 10K-odd neighborhoods, over the whole grid $\mathcal{X}$, compare to the laGP software default of $n=50$. Figure 5 shows how that choice (between the vertical lines) is not advantageous for the Phantom data. Many smaller, but also many much larger larger neighborhoods are estimated by OLAGP, and with much greater frequency. The figure also provides neighborhoods for the new method (“Modular JGP”) discussed next in Section 4.
Before turning to that, we note anecdotally that OLAGP performs worst at or near the manifold of discontinuity (See Appendix A.1). This is because it cannot choose $\hat{n}(x)$ small enough to exclude points from the opposite side of the manifold without maintaining a viable minimum ($n_{\min}=12$) for a GP fit. Observe how the neighborhood for $(0.39,0.34)$ contains several points in the red region. At $\hat{n}(x)=14$ for that location, the neighborhood can’t shrink much further. Consequently, the resulting mean predictions (4) are weighted averages of red- and yellow-region response values. This will yield predictions between 0.4 and 0.6, which we know from Figure 3 are inaccurate. Shrinking neighborhoods farther, even below $n_{\min}=12$, is a double-edged sword: less data means greater reactivity, but also lower information content. It would be better to allow a more flexible notion of neighborhood, beyond simple Euclidean distance in $X_{N}$. Next, we suggest how that can be done without fundamentally altering the LAGP setup by introducing a latent variable. To foreshadow, observe that the “Modular JGP” neighborhood sizes summarized in Figure 5 show fewer small neighborhoods, and fewer large neighborhoods compared to OLAGP.
@@ 4.1 Identifying and predicting levels
Data $Y_{N}$ from a jump process can often be partitioned marginally (i.e., $y$ without $x$), which can inform a more complex spatial partition over $x$. To see this, consider again the histogram of the marginal $Y_{N}$ for the Phantom data [Figure 3, right]. There are two, roughly symmetric, well-separated densities: centered at 0.2 and 0.8. These modes are so well separated you can eyeball it, say at $y=0.5$.
@@
Let $c_{i}\in\{1,2\}$, $i=1,\dots,N$ denote the level learned for each $y_{i}\in Y_{N}$ via model based clustering, with vector $C_{N}$ holding all $N$ levels. For example, we could have $c_{i}=2$ correspond to “yellow” and 1 for “red”, or vice-versa. Then, pair these with inputs $X_{N}$ to form $(X_{N},C_{N})$, comprising “data” for a spatial classification task. Here we shall write $\hat{f}(\cdot)=\mathrm{class}(X_{N},C_{N})$, generically, for the predictive equations of a fitted, spatial classification. One option is (basis-expanded) linear logistic regression, or logistic GP classification (e.g Rasmussen and Williams,, 2006, Chapter 3). Any accurate method will do; our preferred implementation is detailed in Section 5. We may then use $\hat{f}$ to predict component membership at testing locations $\mathcal{X}$: $\hat{f}(x)=\mathbb{P}(y(x)=1)$ for any $x\in\mathcal{X}$. Predicted probabilities for the Phantom data are shown in the left panel of Figure 7.
Figure 7: Left: Jump feature. Center: Nearest 160 neighbors to $x=(-0.32,0.17)$ after adding the jump feature. White points are closest to $x$, and black points are furthest away. Right: Modular JGP fit. Blue dots are two predictive locations and the surrounding black dots are the selected neighborhood.
@@ 4.2 A new feature for GP regression
For this to work, we need to know the probability of component membership for both $Y_{N}$ and $\mathcal{Y}$. Section 4.1 provides estimates of $C_{N}$, for $Y_{N}$, and $\hat{f}(\mathcal{X})$, for $\mathcal{Y}$. The issue is that $C_{N}$ is not a probability; it is a vector consisting of 1s and 2s describing the most likely component a point in $Y_{N}$ belongs to. However, $\hat{f}(\cdot)$ may be used to predict, in-sample, the probability that $Y_{N}$ belongs to a particular component. Vectors $\hat{f}(X_{N})$ and $\hat{f}(\mathcal{X})$ comprise the training and testing jump features for $X_{N}$ and $\mathcal{X}$, respectively. These may be appended as new columns $X\equiv[X,\hat{f}(X)]$ and $\mathcal{X}\equiv[\mathcal{X},\hat{f}(\mathcal{X})]$, overloading the notation somewhat.
@@
To illustrate the effect this new feature has, consider the center plot of Figure 7, again focusing on $x=(-0.32,0.17)$ and its nearest neighbor $x_{*}$. We also plot the 160 NNs to $x$, color coded by the rank of their distance from $x$. The points that are the lightest are closest to $x$, and those that are the darkest are furthest away. Note that these distances are calculated prior to hyperparameter fitting. Observe first that the neighborhood is no longer circular; it has become oval-shaped to avoid the yellow region. Any numbered points that are in the yellow region are much further away than their 2d distance would suggest. Their colors are darker than nearby points the red region, and thus farther from $x$.
@@
We generically notate the fitting stage as $\mathrm{GP}([\mathcal{X},\hat{f}(\mathcal{X})];[X_{N},\hat{f}(X_{N})],Y_{N})$. where “GP” can stand in for any type (e.g., global, LAGP, OLAGP) depending on the problem at hand. We refer to this method, beginning with the clustering algorithm and ending with the final GP fit, as a “Modular JGP”. Algorithm 2 provides the details for quick reference.
Algorithm 2 Outline of Modular JGP
1:$X_{N},Y_{N},\mathcal{X}$
2:$\hat{C}_{N}=\text{clust}(Y_{N})$triangleright perform clustering algorithm on marginal $Y_{N}$, yielding $C_{N}\in\{1,2\}^{N}$
3:$\hat{f}(\cdot)=\text{class}(X_{N},C_{N})$triangleright build classifier mapping $X_{N}$ to $C_{N}$, yielding $\hat{f}(x)=\text{Pr}(y(x)=1)$
4:$\text{GP}([\mathcal{X},\hat{f}(\mathcal{X})];[X_{N},\hat{f}(X_{N})],Y_{N})$triangleright fit any GP (e.g. OLAGP) with jump feature
@@
The right panel of Figure 7 provides a Modular JGP fit on the Phantom data, contrasting with Figure 4. In this case the GP used is an OLAGP, chosen to illustrate the change in neighborhood structure induced by the Modular JGP. Here, in Figure 7, both neighborhoods are smaller, but are completely contained to the same side of the manifold. The Modular JGP has visually improved prediction both visually and via the root mean square error (RMSE; see Appendix B.1) over the global GP, LAGP, and OLAGP.
@@
Now return to Figure 5, summarizing all 10K-odd $\hat{n}(x)$ for OLAGP and Modular JGP. Again, the choice of $n=50$ (the LAGP default) is not particularly advantageous, except for a small percentage of $x\in\mathcal{X}$. Notice that the Modular JGP slightly prefers more mid-range neighborhood sizes than the optimal neighbors GP. In particular, many fewer small $\hat{n}(x)$ are chosen since the neighborhood shape is no longer circular in the original input space (i.e., ignoring the new feature). This facilitates excluding points from the opposite component without resorting to extremely small neighborhood sizes.
@@
Our Modular JGP is purposefully generic: clustering, classifying, predicting, and then fitting with a new feature. We mentioned some of the possible choices for each step (e.g. EM, CGP, and OLAGP), but any methods for clustering, classifying, and regression are fair game. The best choices are likely context/data-dependent. We aim to show that our preferred instance of Modular JGP is an improvement on existing methods in the context of surrogate modeling of jump processes from the recent literature.
@@ 5 Implementation and empirical benchmarking
Here we present three simulated and one real-world example(s) to demonstrate the effectiveness of OLAGP and the Modular JGP. A summary of each data set/experiment is included in Table 1. Details are provided in separate sub-sections to follow. Broadly, for each data set, we conduct a Monte Carlo (MC) experiment which repeatedly randomly partitions the data into 90:10 training and testing sets. In each each repetition, we sample a random subset of 90% of the data set to be used as the training set, and use the remaining 10% as the testing set. The number of MC trials depends on the computational expense implied by the training data set size $n$, which varies in the experiments. We present five competing methods:
@@
  • Original JGP via Park, (2022)’s MATLAB library using classification EM and quadratic boundaries.
  • Modular JGP from Section 4 with further details provided in Table 1.
All experiments were completed in R. Code is provided at https://bitbucket.org/gramacylab/jumpgp. The Original JGP leveraged R.matlab (Bengtsson,, 2022).
Data $N$ $p$ Clustering Classifier GP Type Reps
Phantom 10,201 2 EM CGP OLAGP 100
Star 10,201 2 EM CGP OLAGP 100
Modified Michalewicz 100,000 4 EM Random Forest OLAGP 30
AMHV transport 9,503 4 EM CGP Global GP 100
@@
OLAGP is parallelized over predictive locations $x\in\mathcal{X}$, turning statistical independence into computational exclusivity via foreach (Microsoft and Weston,, 2022). For the Modular JGP, we utilize EM-based clustering and either a CGP (Williams and Barber,, 1998) or a random forest (Breiman,, 2001) for classification via mclust (Scrucca et al.,, 2023), kernlab (Karatzoglou et al.,, 2004) and randomForest (Liaw and Wiener,, 2002), respectively. We only prefer a random forest over a CGP when $n$ is large, for example during the Modified Michalewicz experiment. Again, see Table 1. We measure model performance out-of-sample using $\mathrm{RMSE}=(\frac{1}{m}\sum_{x\in\mathcal{X}}(y(x)-\mu(x))^{2})^{1/2},$ for which lower is better.
@@ 5.1 Phantom
In Sections 34, we assessed model performance (visually and with the RMSE) on the Phantom data for a single train–test partition. The outcome of a full MC experiment is summarized by boxplots in the left panel of Figure 8.
Figure 8: Left: MC experiment on the Phantom data set; Right: A line plot showing the RMSE at individual runs for each of the five methods. The Modular JGP performs best in 96 of 100 runs.
For a more granular view, the right panel draws black lines conneting RMSEs from the same rep, allowing us to see the best performing method for every training/testing set combination. We conclude that most of the overlap of the boxplots is due to variability in the training/testing split. The ordering of accuracy from one MC rep to the next is consistent across methods.
In these plots (and all plots in this section), our two methods are shown third from right (OLAGP) and furthest right (Modular JGP). Both show improvement over the global GP and LAGP, although OLAGP is typically outperformed by the Original JGP. This is unsurprising, since the Original JGP was designed specifically for data from jump processes, and OLAGP was not. But notice that the Modular JGP consistently outperforms the Original JGP, for this data set and broadly over the next few subsections. Looking first at the line plot on the right of Figure 8, we see that for almost every data set (in particular, 96 of 100), predictions made by the Modular JGP are more accurate than those made by the Original JGP. Additionally, the boxplot shows that the Original JGP has higher variability than the Modular JGP.
@@ 5.2 Star
Figure 9: Model fit and chosen neighborhoods for Star data set. Left: Ground truth; Center: OLAGP fit of Star data set; Right: Modular JGP fit of Star data set. Two predictive locations, in blue, are shown with their chosen neighborhoods, in black.
Repeating themes from the Phantom analysis, consider model performance for a single Star train–test split. Mean fits for the global GP and LAGP are shown in Figure 16 and in Appendix B.2, whereas those of OLAGP and the Modular JGP are shown here in the right two panels of Figure 9. RMSEs can be found in Appendix B.2, where we see that the local methods perform well, with the Modular JGP producing the best predictions of the four. The scenarios depicted in right two panels focus on $x=(0.15,0.8)$ and $(0.83,0.11)$. Compared to LAGP, OLAGP has chosen a neighborhood of roughly the same size for $(0.83,0.11)$, and much smaller for $(0.15,0.8)$; however, this smaller neighborhood still contains points from the orange regime. The Modular JGP provides an interesting change: a much smaller neighborhood size for $(0.15,0.11)$, and a larger neighborhood size for $(0.83,0.11)$. Both neighborhoods have adapted their shape to only include points from the same regime, and this improves prediction overall.
Figure 10: Left: MC experiment on the Star data set; Right: A line plot showing the RMSE at individual runs for each of the five methods. The Modular JGP performs best in 76 of 100 runs.
Figure 10 summarizes a MC experiment along similar lines. Here again, OLAGP outperforms both the global GP and LAGP; the two JGP methods typically outperform OLAGP. The Original JGP is again highly variable. Although it outperforms the Modular JGP more frequently than for the Phantom (Original JGP was better in 24 of 100 runs), the Modular JGP is still the best performing method for the Star.

### v2

@@ 3.2 An Illustration
Now consider applying Alg. 1 independently on each $x\in\mathcal{X}$, e.g., testing on a $101\times 101$ grid, extending our running “Phantom” example. The right panel of Figure 4 uses color to indicate the predictive means obtained at each $x$ on the last line of the algorithm. Compared to the truth [Figure 3] the fit looks good, but not perfect. We show in Appendix B.1 that this fit is more accurate by root mean square prediction error (RMSPE) than an ordinary GP fit. Here we aim to illustrate the optimal choice of neighborhood size, $\hat{n}(x)$, in two ways. We begin with focus on two particular predictive locations: (0.32,0.17)(-0.32,0.17) and $(0.39,0.34)$, far from and near to the manifold of discontinuity, respectively. Observe how the optimal neighborhood size (OLAGP) adjusts as appropriate ($n=43$ far from the manifold; $n=14$ near).
Size OLAGP MJGP
12–20 4548 3829
21–40 2838 3013
41–60 745 830
61–80 414 772
81–100 319 277
101–120 139 391
121–140 511 605
141–160 687 484
Figure 5: Distribution of OLAGP and MJGP [Section 4] estimated neighborhood sizes.
Now consider how all 10K-odd neighborhoods, over the whole grid $\mathcal{X}$, compare to the laGP software default of $n=50$. Figure 5 shows how that choice (between the vertical lines) is not advantageous for the Phantom data. Many smaller, but also many much larger larger neighborhoods are estimated by OLAGP, and with much greater frequency. The figure also provides neighborhoods for the new method (“MJGP”) discussed next in Section 4.
Before turning to that, we note anecdotally that OLAGP performs worst at or near the manifold of discontinuity (See Appendix A.1). This is because it cannot choose $\hat{n}(x)$ small enough to exclude points from the opposite side of the manifold without maintaining a viable minimum ($n_{\min}=12$) for a GP fit. Observe how the neighborhood for $(0.39,0.34)$ contains several points in the red region. At $\hat{n}(x)=14$ for that location, the neighborhood can’t shrink much further. Consequently, the resulting mean predictions (4) are weighted averages of red- and yellow-region response values. This will yield predictions between 0.4 and 0.6, which we know from Figure 3 are inaccurate. Shrinking neighborhoods farther, even below $n_{\min}=12$, is a double-edged sword: less data means greater reactivity, but also lower information content. It would be better to allow a more flexible notion of neighborhood, beyond simple Euclidean distance in $X_{N}$. Next, we suggest how that can be done without fundamentally altering the LAGP setup by introducing a latent variable. To foreshadow, observe that the “MJGP” neighborhood sizes summarized in Figure 5 show fewer small neighborhoods, and fewer large neighborhoods compared to OLAGP.
@@ 4.1 Identifying and predicting levels
Data $Y_{N}$ from a jump process can often be partitioned marginally (i.e., $y$ without $x$), which can inform a more complex spatial partition over $x$. To see this, consider again the histogram of the marginal $Y_{N}$ for the Phantom data [Figure 3, right]. There are two, roughly symmetric, well-separated densities, centered at 0.2 and 0.8. These modes are so well separated you can eyeball it, say at $y=0.5$.
@@
Let $c_{i}\in\{1,2\}$, $i=1,\dots,N$ denote the level learned for each $y_{i}\in Y_{N}$ via model based clustering, with vector $C_{N}$ holding all $N$ levels. For example, we could have $c_{i}=2$ correspond to “yellow” and 1 for “red”, or vice-versa. Then, pair these with inputs $X_{N}$ to form $(X_{N},C_{N})$, comprising “data” for a spatial classification task. Here we shall write $\hat{f}(\cdot)=\mathrm{class}(X_{N},C_{N})$, generically, for the predictive equations of a fitted, spatial classification. One option is (basis-expanded) linear logistic regression, or logistic GP classification (e.g Rasmussen and Williams,, 2006, Chapter 3). Any implementation works, although the accuracy of a particular method on a particular data set should be tested before proceeding; our preferred implementation is detailed in Section 5. We may then use $\hat{f}$ to predict component membership at testing locations $\mathcal{X}$: $\hat{f}(x)=\mathbb{P}(y(x)=1)$ for any $x\in\mathcal{X}$. Predicted probabilities for the Phantom data are shown in the left panel of Figure 7.
Figure 7: Left: Jump feature. Center: Nearest 160 neighbors to x=(0.32,0.17)x=(-0.32,0.17) after adding the jump feature. White points are closest to $x$, and black points are furthest away. Right: MJGP fit. Blue dots are two predictive locations and the surrounding black dots are the selected neighborhood.
@@ 4.2 A new feature for GP regression
[…]
@@
To illustrate the effect this new feature has, consider the center plot of Figure 7, again focusing on x=(0.32,0.17)x=(-0.32,0.17) and its nearest neighbor $x_{*}$. We also plot the 160 NNs to $x$, color coded by the rank of their distance from $x$. The points that are the lightest are closest to $x$, and those that are the darkest are furthest away. Note that these distances are calculated prior to hyperparameter fitting. Observe first that the neighborhood is no longer circular; it has become oval-shaped to avoid the yellow region. Any numbered points that are in the yellow region are much further away than their 2d distance would suggest. Their colors are darker than nearby points the red region, and thus farther from $x$.
@@
We generically notate the fitting stage as $\mathrm{GP}([\mathcal{X},\hat{f}(\mathcal{X})];[X_{N},\hat{f}(X_{N})],Y_{N})$, where “GP” can stand in for any type (e.g., global, LAGP, OLAGP) depending on the problem at hand. We refer to this method, beginning with the clustering algorithm and ending with the final GP fit, as a “MJGP”. Algorithm 2 provides the details for quick reference.
Algorithm 2 Outline of MJGP
[…]
@@
The right panel of Figure 7 provides a MJGP fit on the Phantom data, contrasting with Figure 4. In this case the GP used is an OLAGP, chosen to illustrate the change in neighborhood structure induced by the MJGP. Here, in Figure 7, both neighborhoods are smaller, but are completely contained to the same side of the manifold. MJGP has visually improved prediction both visually and via RMSPE (see Appendix B.1) over the global GP, LAGP, and OLAGP.
@@
Now return to Figure 5, summarizing all 10K-odd $\hat{n}(x)$ for OLAGP and MJGP. Again, the choice of $n=50$ (the LAGP default) is not particularly advantageous, except for a small percentage of $x\in\mathcal{X}$. Notice that MJGP slightly prefers more mid-range neighborhood sizes than the optimal neighbors GP. In particular, many fewer small $\hat{n}(x)$ are chosen since the neighborhood shape is no longer circular in the original input space (i.e., ignoring the new feature). This facilitates excluding points from the opposite component without resorting to extremely small neighborhood sizes.
Note that while we have thus far chosen examples where the magnitude of the jump is large compared to the overall variability of the data, that is not required to use MJGP. Data sets with large jumps will see the most improvement when compared to regular GP methods, but using MJGP when jumps are small or nonexistent will not lead to overfitting. See examples of MJGP fit to data with small or nonexistent jumps in Figures 19 and 20 in Appendix C. Although the benefit of using MJGP lessens as the size of the jump decreases, it will not compromise results when jumps are not present.
@@
Our MJGP is purposefully generic: clustering, classifying, predicting, and then fitting with a new feature. We mentioned some of the possible choices for each step (e.g. EM, CGP, and OLAGP), but the choice of methods for clustering, classifying, and regression should be tailored to the context of the problem and the data involved. We aim to show that our preferred instance of MJGP is an improvement on existing methods in the context of surrogate modeling of jump processes from the recent literature.
@@ 5 Implementation and empirical benchmarking
Here we present three simulated and one real-world example(s) to demonstrate the effectiveness of OLAGP and MJGP. A summary of each data set/experiment is included in Table 1; we do not list the clustering algorithm because it did not vary between experiments. Input and response ranges are listed prior to any scaling. Details are provided in separate sub-sections to follow. Broadly, for each data set, we conduct a Monte Carlo (MC) experiment which repeatedly randomly partitions the data into 90:10 training and testing sets. In each each repetition, we sample a random subset of 90% of the data set to be used as the training set, and use the remaining 10% as the testing set. The number of MC trials depends on the computational expense implied by the training data set size $N$, which varies in the experiments. We present five competing methods:
@@
  • Original JGP (OJGP) via Park, (2022)’s MATLAB library using CEM and quadratic boundaries.
  • MJGP from Section 4 with further details provided in Table 1.
All experiments were completed in R. Code is provided at https://bitbucket.org/gramacylab/jumpgp. In this code we provide functions aGP.optneigh.par(), mjgp(), and jgp() to run OLAGP, MJGP, and OJGP, respectively. OJGP leveraged R.matlab (Bengtsson,, 2022).
[…]
Phantom 10,201 2 $0.5,0.5]^{2}$ (17,26)(-17,26) CGP OLAGP 100
Star 10,201 2 $0,1]^{2}$ (16,15)(-16,15) CGP OLAGP 100
Mod. Michalewicz 100,000 4 $0,\pi]^{4}$ (3.25,0.5)(-3.25,0.5) RF OLAGP 30
AMHV Transport 9,503 4
$50,400]\times[1,5]\times$
$0.25,5]\times[0.5,1.5]$
$(100,517)$ CGP Global GP 100
@@
OLAGP is parallelized over predictive locations $x\in\mathcal{X}$, turning statistical independence into computational exclusivity via foreach (Microsoft and Weston,, 2022). For MJGP, we utilize EM-based clustering and either a CGP (Williams and Barber,, 1998) or a random forest (RF; Breiman,, 2001) for classification via mclust (Scrucca et al.,, 2023), kernlab (Karatzoglou et al.,, 2004) and randomForest (Liaw and Wiener,, 2002), respectively. We only prefer a RF over a CGP when $n$ is large, for example during the Modified Michalewicz experiment. Again, see Table 1. Model performance is measured out-of-sample using $\mathrm{RMSPE}=(\frac{1}{m}\sum_{x\in\mathcal{X}}(y(x)-\mu(x))^{2})^{1/2},$ for which lower is better. Computation times are in Appendix D and prediction interval coverage is in Appendix E.
@@ 5.1 Phantom
In Sections 34, we assessed model performance (visually and with RMSPE) on the Phantom data for a single train–test partition. The outcome of a full MC experiment is summarized by boxplots in the left panel of Figure 8.
Figure 8: Left: MC experiment on the Phantom data set; Right: A line plot showing RMSPE at individual runs for each of the five methods. MJGP performs best in all runs.
For a more granular view, the right panel draws black lines connecting RMSPEs from the same rep, allowing us to see the best performing method for every training/testing set combination. We conclude that most of the overlap of the boxplots is due to variability in the training/testing split. The ordering of accuracy from one MC rep to the next is consistent across methods.
In these plots (and all plots in this section), our two methods are shown third from right (OLAGP) and furthest right (MJGP). Both show improvement over the global GP and LAGP, although OLAGP is typically outperformed by OJGP. This is unsurprising, since OJGP was designed specifically for data from jump processes, and OLAGP was not. But notice that MJGP consistently outperforms OJGP, here and broadly over the next few subsections. Looking first at the line plot on the right of Figure 8, we see that for every data set split, predictions made by MJGP are more accurate than those from OJGP. Additionally, the boxplot shows that OJGP has higher variability than MJGP. This added variability is explained by the use of a quadratic rather than linear boundary; for more on this distinction, see Park, (2022).
@@ 5.2 Star
Figure 9: Model fit and chosen neighborhoods for Star data set. Left: Ground truth; Center: OLAGP fit of Star data set; Right: MJGP fit of Star data set. Two predictive locations, in blue, are shown with their chosen neighborhoods, in black.
Repeating themes from the Phantom analysis, consider model performance for a single Star train–test split. Mean fits for the global GP and LAGP are shown in Figure 17 and in Appendix B.2, whereas those of OLAGP and MJGP are shown here in the right two panels of Figure 9. RMSPEs can be found in Appendix B.2, where we see that the local methods perform well, with MJGP producing the best predictions of the four. The scenarios depicted in right two panels focus on $x=(0.15,0.8)$ and $(0.83,0.11)$. Compared to LAGP, OLAGP has chosen a neighborhood of roughly the same size for $(0.83,0.11)$, and much smaller for $(0.15,0.8)$; however, this smaller neighborhood still contains points from the orange regime. MJGP provides an interesting change: a much smaller neighborhood size for $(0.15,0.11)$, and a larger neighborhood size for $(0.83,0.11)$. Both neighborhoods have adapted their shape to only include points from the same regime, and this improves prediction overall.
Figure 10: Left: MC experiment on the Star data set; Right: A line plot showing RMSPE at individual runs for each of the five methods. MJGP performs best in 82 of 100 runs.
Figure 10 summarizes a MC experiment along similar lines. Here again, OLAGP outperforms both the global GP and LAGP; the two JGP methods typically outperform OLAGP. OJGP is again highly variable. Although it outperforms MJGP more frequently than for the Phantom (OJGP was better in 17 of 100 runs), MJGP is still the best performing method for the Star.